# Soft decision trees for survival analysis

### v1

@@ Appendix B Survival tree performance measures
In this appendix, we briefly recall the accuracy measures considered in this work to evaluate the accuracy of our soft survival trees which are commonly used to assess the performance of survival models. Performance measures for survival models generally focus on their discrimination and calibration capabilities. Discrimination refers to a model’s ability to distinguish between subjects who experience an event and those who do not, assigning higher predicted probabilities to the former. Calibration assesses how closely the predicted probabilities match the observed outcomes. The reader is referred to (Steyerberg, 2009) for a detailed description.
@@ B.0.1 Concordance Indices
The Concordance Statistic, a well-established measure of discrimination, has been extended from logistic regression to survival models Harrell et al. (1982). Harrell’s C-index is a concordance statistic based on the principle that a good survival model should assign lower survival probabilities to data points with shorter times-to-event.
@@
A main drawback of Harrell’s C is its bias in datasets with a high proportion of censored data. To mitigate this limitation, Uno et al. (2011) enhanced Harrell’s C by applying Inverse Probability of Censoring Weights (IPCW), making it more robust to extensive censoring. The Uno’s C-index is given by:
@@ B.0.2 Cumulative Dynamic AUC
where $\hat{G}(\cdot)$ is the usual IPCW computed using the Kaplan-Meier estimator.
Lambert and Chevret (2016) proposed the Cumulative Dynamic AUC (CD-AUC), which restricts the time-dependent weighted AUC estimator to a specific time interval, providing a summary measure of the mean AUC over the selected range:
@@
where $\hat{S}(\cdot)$ is the Kaplan-Meier estimator of the survival function and the time range is derived from the dataset as $t_{\text{min}}=\min\{t_{i}\}_{i=1}^{N}$ and $t_{\text{max}}=\max\{t_{i}\}_{i=1}^{N}$.
@@ B.0.3 Integrated Brier score
The Brier Score (BS) is a widely used measure for assessing the accuracy of probabilistic classification models. Graf et al. (1999) introduced an adapted version of the BS to handle survival datasets with censored outputs. In survival analysis, BS is commonly used to evaluate the calibration of survival models. At a specific time $t$, $BS(t)$ represents the mean squared error between the observed data and the predicted survival function $\hat{S}_{\mathbf{x}_{i}}(t)$, adjusted using IPCW:
@@
Similar to the cdAUC, the Integrated Brier Score (IBS) offers a comprehensive assessment of model performance across a specified time interval:
[…]
@@ C.0.1 Initialization procedure
Since the NLO training formulation (5) is nonconvex, appropriate initial solutions can improve the training phase and enhance testing accuracy. To find suitable initial solutions, we adapt the initialization procedure proposed in (Consolo et al., 2025) for soft regression trees to survival analysis. Starting from the root node with all the input vectors $\mathbf{x}_{i}$ of the training set $I$, we proceed level by level towards the leaf nodes and we apply at each branch node $n\in\tau_{B}$ a binary clustering algorithm in order to partition the input vectors assigned to node $n$ into two subsets of input vectors to be assigned to the two child nodes of $n$, denoted by $X^{right}_{n}$ and $X^{left}_{n}$. Once each input vector has been assigned to a single root-to-leaf-node path, we determine for each branch node $n$ the corresponding values of the variables $\bm{\omega}^{0}_{n}$ by solving a binary Logistic Regression problem where we try to route all input vectors in $X^{right}_{n}$ ($X^{left}_{n}$) along the right (left) branch of node $n$. Similarly, the corresponding variables $\bm{\beta}^{0}_{n}$ characterizing the survival curves associated to each leaf node $n$ are determined by solving a maximum likelihood problem where only the input vectors falling into the leaf node $n$ are considered. It is worth pointing out that this initialization procedure is applicable to any type of (parametric or semiparametric) survival curves which can be estimated via a maximum likelihood approach.
This sequence of binary clustering problems from the root node to the bottom level leaf nodes is repeated several times. At the end, the set of clusters associated to the leaf nodes that maximizes an appropriate clustering quality metric is used to generate an initial solution $(\bm{\omega}^{0},\bm{\beta}^{0})$. The quality of the clusters at the leaf nodes is assessed using silhouette scores, which indicate the average similarity between each cluster and its most similar one (Dudek, 2020). As in Consolo et al. (2025), we adopt the Davies-Bouldin index as the silhouette score (Davies and Bouldin, 1979), which represents the average similarity of each cluster to its most similar cluster. Once the best set of clusters of input vectors associated to the leaf nodes is selected based on the silhouette score, the variables $\bm{\omega}^{0}$ and $\bm{\beta}^{0}$ for all the nodes are determined as above. The reader is referred to (Consolo et al., 2025) for details on the initialization procedure.
@@ C.0.2 Branch node update procedure
1:procedure UpdateBranchNode($I$,$W^{k}_{B},\bm{\omega}^{k}_{W^{k}_{B}},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}$)
2:     $s\leftarrow$ $\min\{{\hat{s}}\ |\;{\hat{s}}\in W^{k}_{B}\}$triangleright s is the ancestor of all the other nodes in $W^{k}_{B}$
3:     $I_{s}\leftarrow$$\{(\textbf{x}_{i},t_{i},c_{i})\in I|$ HBP path of $(\textbf{x}_{i},t_{i},c_{i})$ contains $s\}$triangleright data points of $I_{s}$ falling into $s$
4:     if ($L_{s}\leq\varepsilon_{1}\;\text{or}\;L_{s}\geq 1-\varepsilon_{1})\;\text{and}\;({\varepsilon_{1}}N\geq 1)$  $\triangleright$ imbalanced data points routing
@@
9:         if $L_{s}\leq\varepsilon_{2}\;\text{or}\;L_{s}\geq 1-{\varepsilon_{2}}$ $\triangleright$ high imbalanced routing
10:              $\bullet$ let $d_{max}\;\text{be the index of the child node of $s$ with maximum number of $\mathbf{x}_{i}$ routed to it}$
@@
17:         $\tilde{\bm{\omega}}_{W^{k}_{B}}\leftarrow\{\tilde{\bm{\omega}}_{s},\tilde{\bm{\omega}}_{W^{k}_{B}\setminus s}\}$triangleright $\tilde{\bm{\omega}}_{s}$ are the only updated variables
@@
19:         determine $\tilde{\bm{\omega}}_{W^{k}_{B}}$ by solving (5) with respect to variables ${\bm{\omega}}_{W^{k}_{B}}$triangleright minimize with any algorithm
@@ Appendix E Interpretability results
Figure 10 presents the survival functions distribution from a single run (a single fold and initial solution) on the Aids__\__death, Aids, Churn, Credit, Gbsg2, Uissurv, and Whas500 datasets. The results for depth $D=1$ are shown in the first column, while those for depth $D=2$ are displayed in the second column. Colors are used to identify specific leaf nodes, with varying shades within each node illustrating differences in the survival functions among data points.
@@
D=1 D=2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Figure 10: The survival function distributions for a single run across the Aids__\__death, Aids, Churn, Credit, Gbsg2, and Whas500 datasets. The results for depth $D=1$ are shown on the left, while those for $D=2$ are on the right. Each color corresponds to a specific leaf node, with variations in shading within each node illustrating the differences in survival functions among individual data points.
@@ Appendix F Fairness results
In this appendix we evaluate the performance of our fairness-promoting formulation (4) optimized using the node-based decomposition training algorithm on a socio-economic dataset (Romeo, 1999) where a sensitive group is identifiable. Moreover, we report some plots to illustrate the impact on the survival function of the penalty term on the individual survival functions as $\rho$ grows.
The $\rho$ hyperparameter begins at 00 and increases to $\frac{20}{N_{M}N_{F}}$ in five steps ($\frac{1}{N_{M}N_{F}}[0,1,5,10,15,20]$), where $N_{M}$ and $N_{F}$ denote the number of males and females in the training dataset.

### v3

@@ Appendix B Survival tree performance measures
In this appendix, we briefly recall the accuracy measures considered in this work to evaluate the accuracy of our soft survival trees which are commonly used to assess the performance of survival models. Performance measures for survival models generally focus on their discrimination and calibration capabilities. Discrimination refers to a model’s ability to distinguish between subjects who experience an event and those who do not, assigning higher predicted probabilities to the former. Calibration assesses how closely the predicted probabilities match the observed outcomes. The reader is referred to (steyerberg2009clinical) for a detailed description.
@@ B.0.1 Concordance Indices
The Concordance Statistic, a well-established measure of discrimination, has been extended from logistic regression to survival models harrell1982evaluating. Harrell’s C-index is a concordance statistic based on the principle that a good survival model should assign lower survival probabilities to data points with shorter times-to-event.
@@
A main drawback of Harrell’s C is its bias in datasets with a high proportion of censored data. To mitigate this limitation, uno2011c enhanced Harrell’s C by applying Inverse Probability of Censoring Weights (IPCW), making it more robust to extensive censoring. The Uno’s C-index is given by:
@@ B.0.2 Cumulative Dynamic AUC
lambert2016summary proposed the Cumulative Dynamic AUC (CD-AUC), which restricts the time-dependent weighted AUC estimator to a specific time interval, providing a summary measure of the mean AUC over the selected range:
$\text{CD-AUC}=\frac{1}{\hat{S}(t_{\text{min}})-\hat{S}(t_{\text{max}})}\int_{t_{\text{min}}}^{t_{\text{max}}}\hat{\text{AUC}}(t)\,d\hat{S}(t)$
[…]
@@ B.0.3 Integrated Brier score
The Brier Score (BS) is a widely used measure for assessing the accuracy of probabilistic classification models. graf1999assessment introduced an adapted version of the BS to handle survival datasets with censored outputs. In survival analysis, BS is commonly used to evaluate the calibration of survival models. At a specific time $t$, $BS(t)$ represents the mean squared error between the observed data and the predicted survival function $\hat{S}_{\mathbf{x}_{i}}(t)$, adjusted using IPCW:
@@
Similar to the CD-AUC, the Integrated Brier Score (IBS) offers a comprehensive assessment of model performance across a specified time interval:
$\text{IBS}=\frac{1}{t_{\text{max}}}\int_{0}^{t_{\text{max}}}BS(t)\,dt$ (15)
@@ C.0.1 Initialization procedure
Since the NLO training formulation (13) is nonconvex, appropriate initial solutions can improve the training phase and enhance testing accuracy. To find suitable initial solutions, we adapt the initialization procedure proposed in (consolo2025) for soft regression trees to survival analysis. Starting from the root node with all the input vectors $\mathbf{x}_{i}$ of the training set $I$, we proceed level by level towards the leaf nodes and we apply at each branch node $n\in\tau_{B}$ a binary clustering algorithm in order to partition the input vectors assigned to node $n$ into two subsets of input vectors to be assigned to the two child nodes of $n$, denoted by $X^{right}_{n}$ and $X^{left}_{n}$. Once each input vector has been assigned to a single root-to-leaf-node path, we determine for each branch node $n$ the corresponding values of the variables $\bm{\omega}^{0}_{n}$ by solving a binary Logistic Regression problem where we try to route all input vectors in $X^{right}_{n}$ ($X^{left}_{n}$) along the right (left) branch of node $n$. Similarly, the corresponding variables $\bm{\beta}^{0}_{n}$ characterizing the survival curves associated to each leaf node $n$ are determined by solving a maximum likelihood problem where only the input vectors falling into the leaf node $n$ are considered. It is worth pointing out that this initialization procedure is applicable to any type of (parametric or semiparametric) survival curves which can be estimated via a maximum likelihood approach.
This sequence of binary clustering problems from the root node to the bottom level leaf nodes is repeated several times. At the end, the set of clusters associated to the leaf nodes that maximizes an appropriate clustering quality metric is used to generate an initial solution $(\bm{\omega}^{0},\bm{\beta}^{0})$. The quality of the clusters at the leaf nodes is assessed using silhouette scores, which indicate the average similarity between each cluster and its most similar one (dudek2020silhouette). As in consolo2025, we adopt the Davies-Bouldin index as the silhouette score (davies1979cluster), which represents the average similarity of each cluster to its most similar cluster. Once the best set of clusters of input vectors associated to the leaf nodes is selected based on the silhouette score, the variables $\bm{\omega}^{0}$ and $\bm{\beta}^{0}$ for all the nodes are determined as above. The reader is referred to (consolo2025) for details on the initialization procedure.
@@ C.0.2 Branch node update procedure
$\textbf{Input:}\ \text{training set}\ I;\ W^{k}_{B}\,\text{working}\,\text{set};\,\text{previous}\;\text{update}\,\bm{\omega}^{k}_{W^{k}_{B}};\,\varepsilon_{1},\,\varepsilon_{2},\,\varepsilon_{3}\in(0,1),\,\varepsilon_{1}>\varepsilon_{2}\\$ $\textbf{Output:}\,\text{update}\,\tilde{\bm{\omega}}_{W^{k}_{B}}$
@@
2:  $s\leftarrow$ $\min\{{\hat{s}}\ |\;{\hat{s}}\in W^{k}_{B}\}$ $\triangleright$ s is the ancestor of all the other nodes in $W^{k}_{B}$
3:  $I_{s}\leftarrow$ $\{(\textbf{x}_{i},t_{i},c_{i})\in I|$ HBP path of $(\textbf{x}_{i},t_{i},c_{i})$ contains s}s\} $\triangleright$ data points of $I_{s}$ falling into $s$
[…]
@@
[…]
@@
17:   $\tilde{\bm{\omega}}_{W^{k}_{B}}\leftarrow\{\tilde{\bm{\omega}}_{s},\tilde{\bm{\omega}}_{W^{k}_{B}\setminus s}\}$ $\triangleright$ $\tilde{\bm{\omega}}_{s}$ are the only updated variables
@@
19:   determine $\tilde{\bm{\omega}}_{W^{k}_{B}}$ by solving (13) with respect to variables ${\bm{\omega}}_{W^{k}_{B}}$ $\triangleright$ minimize with any algorithm
@@ Appendix E Interpretability results
Figure 10 presents the survival functions distribution from a single run (a single fold and initial solution) on the Aids_\_death, Aids, Churn, Credit, Gbsg2, Uissurv, and Whas500 datasets. The results for depth $D=1$ are shown in the first column, while those for depth $D=2$ are displayed in the second column. Colors are used to identify specific leaf nodes, with varying shades within each node illustrating differences in the survival functions among data points.
@@
D=1 D=2 [Uncaptioned image] [Uncaptioned image]
@@
Figure 10: The survival function distributions for a single run across the Aids_\_death, Aids, Churn, Credit, Gbsg2, and Whas500 datasets. The results for depth $D=1$ are shown on the left, while those for $D=2$ are on the right. Each color corresponds to a specific leaf node, with variations in shading within each node illustrating the differences in survival functions among individual data points.
@@ Appendix F Fairness results
In this appendix we evaluate the performance of our fairness-promoting formulation (12) optimized using the node-based decomposition training algorithm on a socio-economic dataset (romeo1999conducting) where a sensitive group is identifiable. Moreover, we report some plots to illustrate the impact on the survival function of the penalty term on the individual survival functions as $\rho$ grows.
The $\rho$ hyperparameter begins at 0 and increases to $\frac{20}{N_{M}N_{F}}$ in five steps ($\frac{1}{N_{M}N_{F}}[0,1,5,10,15,20]$), where $N_{M}$ and $N_{F}$ denote the number of males and females in the training dataset.